# Computation and Communication Co-scheduling for Timely Multi-Task Inference at the Wireless Edge

### v1

@@ IV-A Lagrangian Relaxation and Dual Problem
To develop an asymptotically optimal policy for (III-A)-(6), following recent techniques for weakly coupled MDPs [32, 12], we first relax the problem using Lagrange multipliers. We associate a vector of non-negative Lagrange multipliers $\boldsymbol{\lambda}_{t}=(\lambda_{1,t},\lambda_{2,t},\ldots,\lambda_{M,t})$ with constraints (5) and a non-negative Lagrange multiplier $\mu_{t}$ with constraint (6) at each time $t$. To avoid an infinite number of Lagrange multipliers associated with the constraints over the infinite time horizon from $t=0$ to $t=\infty$, we truncate the problem to a finite time horizon $T$ as shown in (IV-A). Due to bounded inference error function, the performance loss resulting from this truncation becomes negligible for sufficiently large values of $T$.
The truncated problem is given by
$\bar{p}(\boldsymbol{\lambda},\boldsymbol{\mu};\tau:T)=$
$\;\inf_{\pi\in\Pi}\sum_{t=\tau}^{T}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\frac{\gamma^{t}\mathbb{E}_{\pi}\left[w_{m,j}p_{m,j}(\Delta_{m,j}(t))\right]}{K}$
$+\sum_{t=\tau}^{T}\sum_{m=1}^{M}{\lambda_{m,t}}\frac{\gamma^{t}}{K}\left(\left(\sum_{j=1}^{k_{m}}\pi_{m,j}(t)\right)-C_{m}\right)$
$+\sum_{t=\tau}^{T}{\mu_{t}}\frac{\gamma^{t}}{K}\left(\left(\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)n_{m,j}\right)-N\right),$ (7)
where $\boldsymbol{\lambda}=(\boldsymbol{\lambda}_{\tau},\boldsymbol{\lambda}_{\tau+1},\ldots,\boldsymbol{\lambda}_{T})$, $\boldsymbol{\mu}=(\mu_{0},\mu_{1},\ldots,\mu_{T})$, $\bar{p}(\boldsymbol{\lambda},\boldsymbol{\mu};\tau:T)$ is the optimal value of (IV-A), and $\tau=0,1,\ldots,T$.
@@
$(\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*})=\operatorname*{arg\,max}_{(\boldsymbol{\lambda},\boldsymbol{\mu})\geq 0}\bar{p}(\boldsymbol{\lambda},\boldsymbol{\mu};\tau:T),$ (8)
where $(\boldsymbol{\lambda}^{*},\boldsymbol{\mu}^{*})$ is the optimal dual solution.
@@ IV-B Optimal Solution to (IV-A)
$\bar{p}_{m,j}(\boldsymbol{\lambda}_{m},\boldsymbol{\mu};\tau:T)$
$=$ $\inf_{\pi_{m,j}\in\Pi_{m,j}}\sum_{t=\tau}^{T}\gamma^{t}\mathbb{E}_{\pi_{m,j}}\bigg{[}w_{m,j}p_{m,j}(\Delta_{m,j}(t))$
@@
where $\bar{p}_{m,j}(\boldsymbol{\lambda}_{m},\mu;\tau:T)$ is the optimal objective value of the sub-problem, $\boldsymbol{\lambda}_{m}=(\lambda_{m,\tau},\lambda_{m,\tau+1},\ldots,\lambda_{m,T}),$ $\pi_{m,j}=(\pi_{m,j}(\tau),\ldots,\pi_{m,j}(T))$ is a scheduling policy for the task, and $\Pi_{m,j}$ is the set of all causal signal-ignorant policies. The Lagrange multipliers $\boldsymbol{\lambda}_{m}$ and $\boldsymbol{\mu}$ correspond to the computation cost and the communication cost terms, respectively.
@@
There exists an optimal policy for (IV-B) in which the optimal decision $\pi^{*}_{m,j}(t)$ at each time $t$ is determined by
$\pi^{*}_{m,j}(t)=\operatorname*{arg\,min}_{\pi_{m,j}(t)\in\{0,1\}}Q^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\Delta_{m,j}(t),\pi_{m,j}(t)),$ (10)
where the action value function $Q^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\cdot,\cdot)$ is given by
$Q^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta,a)$ $=w_{m,j}p_{m,j}(\delta)+(1-a)\gamma V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t+1}(\delta+1)$
$+a\left(\lambda_{m,t}+\mu_{t}n_{m,j}+\gamma V^{\boldsymbol{\lambda_{m}},\boldsymbol{\mu}}_{m,j,t+1}(1)\right),$ (11)
the value function $V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\cdot)$ for all $\delta\in\mathbb{Z}^{+}$ and $t=\tau,\tau+1,\ldots,T$ is given by
$V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta)=\min_{a\in\{0,1\}}Q^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta,a),$ (12)
@@
$V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,T+1}(\delta)=0.$ (13)
@@
$\!\!\!\!\!V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta)$
@@
$\!\!\!\!\!+\gamma\mathbb{E}\!\left[V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t+1}(\Delta_{m,j}(t+1))\bigg{|}\Delta_{m,j}(t)\!=\!\delta,\!\pi_{m,j}(t)\!=a\!\right].$ (14)
Lemma 1 establishes an optimal decision $\pi^{*}_{m,j}(t)$ for problem (IV-B) by using dynamic programming method. The backward induction method to compute the value function $V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta)$ for $t=\tau,\tau+1,\ldots,T$ is given by
$V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta)\!$ $=\!w_{m,j}p_{m,j}(\delta)+\!\min_{a\in\{0,1\}}\bigg{\{}\!(1-a)\gamma V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t+1}(\delta+1)$
$+a\left(\lambda_{m,t}+\mu_{t}n_{m,j}+\gamma V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t+1}(1)\right)\bigg{\}}.$ (15)
However, if the AoI $\delta$ can take infinite values, this is computationally intractable. Thus, we restrict the computation of value function to a finite range $\delta=1,2,\ldots,\bar{\delta}$, and approximate $V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\delta)\approx V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\bar{\delta})$ for values exceeding this range. In reality, this truncation will have a negligible effect since (i) higher AoI values are rarely visited in practice [37], and (ii) the inference error $p_{m,j}(\delta)$ tends to converge to an upper bound as AoI becomes large, as seen in some recent works [3, 9, 10] and our machine learning experiments in Fig. 4. The backward induction algorithm has a time complexity of $O(\bar{\delta}T)$.
@@ IV-C Solution to (8)
Next, we solve the dual problem (8). We can reformulate the dual problem as
$\max_{\boldsymbol{\lambda}\geq 0,\boldsymbol{\mu}\geq 0}$ $\sum_{t=\tau}^{T}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\Delta_{m,j}(t))$
$-\sum_{t=\tau}^{T}\sum_{m=1}^{M}\gamma^{t-\tau}\lambda_{m,t}C_{m}+\sum_{t=\tau}^{T}\gamma^{t-\tau}\mu_{t}N,$ (16)
where $V^{\boldsymbol{\lambda}_{m},\boldsymbol{\mu}}_{m,j,t}(\Delta_{m,j}(t))$ is obtained by using (IV-B).
@@ V Scheduling Policy
V-A Maximum Gain First (MGF) Scheduler
While the decision $\pi^{*}_{m,j}(t)$ provided in Lemma 1 may violate constraints (5)-(6), we exploit the structure of the decision $\pi^{*}_{m,j}(t)$ to develop a scheduling policy for the original problem (III-A)-(6). The proposed policy utilizes the notion of gain indices discussed in some recent papers [3, 19, 27]. To determine gain indices for our MTRI problem at time $t$, we use the action value function $Q^{\boldsymbol{\lambda}^{*}_{m},\boldsymbol{\mu}^{*}}_{m,j,t}(\cdot)$ associated with Lagrange multipliers $\lambda^{*}_{m,t},\lambda^{*}_{m,t+1},\ldots,\lambda^{*}_{m,T}$ and $\mu^{*}_{t},\mu^{*}_{t+1},\ldots,\mu^{*}_{T}$.
@@ V-A Maximum Gain First (MGF) Scheduler
$\alpha_{m,j,t}(\delta)=Q^{\boldsymbol{\lambda}^{*}_{m},\boldsymbol{\mu}^{*}}_{m,j,t}(\delta,0)-Q^{\boldsymbol{\lambda}^{*}_{m},\boldsymbol{\mu}^{*}}_{m,j,t}(\delta,1),$ (17)
where $(\boldsymbol{\lambda}^{*}_{m},\boldsymbol{\mu}^{*})$ is an optimal solution to (IV-A) after substituting $\tau=t$ in (IV-A).
Algorithm 1 presents our maximum gain first (MGF) scheduler for solving the main problem (III-A)-(6). At each time $t$, the policy $\pi_{\text{MGF}}$ prioritizes generating and transmitting features ($\pi_{m,j}(t)=1$) for the inference tasks with highest gain index, while adhering to the available communication and computation resources. Formally, at time $t$, let
$A(t)=\bigg{\{}(m,j):\alpha_{m,j,t}(\Delta_{m,j}(t))>0\bigg{\}}$ (18)
be the set of inference tasks with positive gain indices. Our policy then proceeds as follows:
  • (1)
    Select the inference task $(m^{*},j^{*})$ that satisfies
  • (2)
V-B Performance Analysis
We now analyze the performance of our policy relative to the original problem (III-A)-(6). Following standard practice in the weakly-coupled MDP literature [12, 30], we consider a set of sub-problems at source $m$ to be in the same class if they share identical penalty functions, weights, and transition probabilities, where each sub-problem is indexed by the $(m,j)$-th inference task.
Definition 2 (Asymptotic optimality).
Consider a “base” MTRI system with $N$ channels, $M$ sources, $k_{m}$ classes of sub-problems per source $m$, and a computation resource budget $C_{m}$ for source $m$. Let $\bar{p}^{r}_{{\text{MGF}}}$ represent the discounted infinite horizon sum of inference errors under policy $\pi_{\text{MGF}}$ for a system with computation resource budget $rC_{m}$, communication resource budget $rN$, and $k_{m}$ classes of sub-problems per source $m$ in which each class of sub-problems contain $r$ inference tasks while maintaining a constant $M$ sources. The policy $\pi_{\text{MGF}}$ is asymptotically optimal if $\bar{p}^{r}_{{\text{MGF}}}=\bar{p}_{{\pi}^{r}_{opt}}$ for all $\pi\in\Pi$ as inference task per class $r$ approaches $\infty$.
First, we provide Lemma 2, which is a key tool to showing asymptotic optimality. Let the policy $\pi^{*}=(\pi^{*}_{m,j}(t))_{\forall m,j,t\leq T}$ be an optimal solution to (IV-A), where the action $\pi^{*}_{m,j}(t)$ is obtained by using Lemma 1 and by getting the optimal Lagrange multipliers of (8) with $\tau=t$. Using (10) and (17), we can verify that $\pi^{*}_{m,j}(t)=1$ if and only if $\alpha_{m,j,t}(\Delta_{m,j}(t))>0$.
Lemma 2.
For any time $t$ and AoI values, the expected number of inference tasks $(m,j)$ with different actions under the MGF policy and the policy $\pi^{*}$ is bounded above by $\sum_{m=1}^{M}\sqrt{k_{m}}+\sqrt{\sum_{m=1}^{M}k_{m}}$.
Proof.
See Appendix A. ∎
Then, we can obtain our main theoretical result:
Theorem 1.
If the weighted functions $w_{m,j}p_{m,j}(\delta)$ are bounded for all tasks $(m,j)$ and
$T\geq\log_{\frac{1}{\gamma}}\left(\sum_{m=1}^{M}\sqrt{rk_{m}}\right),$ (20)
then the MGF policy is asymptotically optimal as the number of inference tasks $r$ per class of sub-problems increases. Specifically, we have
$=O\bigg{(}\frac{1}{\sum_{m=1}^{M}\sqrt{rk_{m}}}\bigg{)},$ (21)
where $k_{m}$ is the number of sub-problems per source, $M$ is the number of sources, $\gamma$ is the discount factor, and $\bar{p}_{h}$ and $\bar{p}_{l}$ are finite constants such that $\bar{p}_{l}\leq w_{m,j}p_{m,j}(\delta)\leq\bar{p}_{h}$ for any sub-problem $(m,j)$ and AoI value $\delta=1,2,\ldots.$
Proof.
See Appendix B. ∎
While prior work has introduced gain-index-based policies for RMAB problems [19, 27, 10], these cannot be directly applied to general weakly-coupled MDPs. Our gain-index-based policy, a specialized re-optimized fluid policy (Definition 3) [12], achieves tighter asymptotic optimality for MTRI systems (see (1)) compared to the $O(\frac{1}{\sqrt{\sum_{m=1}^{M}rk_{m}}})$ bound in [12]. This improvement is obtained by using Lemma 2, which strengthens the result [12, Lemma EC1.1] by exploiting the MTRI constraint structure: sub-problems utilize only their source’s computational resources. Unlike the general system in [12] where all resources are globally shared, MTRI systems have local (computational) and global (communication) resources, yielding a tighter bound. For example, a sub-problem associated with source $m_{1}$ would only consume computational resources from source $m_{1}$ and not from another source $m_{2}$, unlike in [12], where all sub-problems share all resources in the system.

### v2

@@ IV-A Lagrangian Relaxation and Dual Problem
To develop an asymptotically optimal policy for (4)-(6), following recent techniques for weakly coupled MDPs [47, 13], we first relax the problem using Lagrange multipliers. We associate a vector of non-negative Lagrange multipliers $\boldsymbol{\lambda}_{t}=(\lambda_{1,t},\lambda_{2,t},\ldots,\lambda_{M,t})$ with constraints (5) and a non-negative Lagrange multiplier $\mu_{t}$ with constraint (6) at each time $t$. To avoid an infinite number of Lagrange multipliers associated with the constraints over the infinite time horizon from $t=0$ to $t=\infty$, we truncate the problem to a finite time horizon $T$ as shown in (IV-A). Due to bounded inference error function, the performance loss resulting from this truncation becomes negligible for sufficiently large values of $T$.
The truncated problem from any time $\tau\in\{0,1,\ldots,T\}$ to $T$ is given by
$\bar{p}_{\tau}(\boldsymbol{\lambda}_{\tau:T},\boldsymbol{\mu}_{\tau:T})=$
$\;\inf_{\pi\in\Pi}\sum_{t=\tau}^{T}\frac{\gamma^{t-\tau}}{K}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\mathbb{E}_{\pi}\left[w_{m,j}p_{m,j}(\Delta_{m,j}(t))\right]$
$+\sum_{t=\tau}^{T}\sum_{m=1}^{M}{\lambda_{m,t}}\frac{\gamma^{t-\tau}}{K}\left(\left(\sum_{j=1}^{k_{m}}\pi_{m,j}(t)\right)-C_{m}\right)$
$+\sum_{t=\tau}^{T}{\mu_{t}}\frac{\gamma^{t-\tau}}{K}\left(\left(\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)n_{m,j}\right)-N\right),$ (7)
where
$\boldsymbol{\lambda}_{t}=$ $(\lambda_{1,t},\lambda_{2,t},\ldots,\lambda_{M,t}),$ (8)
$\boldsymbol{\lambda}_{\tau:T}=$ $(\boldsymbol{\lambda}_{\tau},\boldsymbol{\lambda}_{\tau+1},\ldots,\boldsymbol{\lambda}_{T}),$ (9)
$\boldsymbol{\mu}_{\tau:T}=$ $(\mu_{\tau},\mu_{\tau+1},\ldots,\mu_{T}),$ (10)
and $\bar{p}_{\tau}(\boldsymbol{\lambda}_{\tau:T},\boldsymbol{\mu}_{\tau:T})$ is the optimal value of (IV-A).
@@
$(\boldsymbol{\lambda}^{*}_{\tau:T},\boldsymbol{\mu}^{*}_{\tau:T})=\operatorname*{arg\,max}_{(\boldsymbol{\lambda}_{\tau:T},\boldsymbol{\mu}_{\tau:T})\geq 0}~\bar{p}_{\tau}(\boldsymbol{\lambda}_{\tau:T},\boldsymbol{\mu}_{\tau:T}),$ (11)
where $(\boldsymbol{\lambda}^{*}_{\tau:T},\boldsymbol{\mu}^{*}_{\tau:T})$ is the optimal dual solution.
@@ IV-B Optimal Solution to (IV-A)
$\inf_{\pi_{m,j}\in\Pi_{m,j}}\sum_{t=\tau}^{T}\gamma^{t-\tau}$ $\mathbb{E}_{\pi_{m,j}}\bigg{[}w_{m,j}p_{m,j}(\Delta_{m,j}(t))$
@@
where $\pi_{m,j}=(\pi_{m,j}(\tau),\ldots,\pi_{m,j}(T))$ is a scheduling policy for the task and $\Pi_{m,j}$ is the set of all causal signal-ignorant policies.
@@
There exists an optimal policy for (IV-B) in which the optimal decision $\pi^{*}_{m,j}(t)$ at each time $t$ minimizes the action value function
$\min_{\pi_{m,j}(t)\in\{0,1\}}Q^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\Delta_{m,j}(t),\pi_{m,j}(t)),$ (13)
where the action value function $Q^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\cdot,\cdot)$ is given by
$Q^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta,a)$
$=w_{m,j}p_{m,j}(\delta)+(1-a)\gamma V^{\boldsymbol{\lambda}_{m,t+1:T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t+1}(\delta+1)$
$\quad+a\left(\lambda_{m,t}+\mu_{t}n_{m,j}+\gamma V^{\boldsymbol{\lambda}_{m,t+1:T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t+1}(1)\right),$ (14)
the value function $V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t}(\delta)$ for all $\delta\in\mathbb{Z}^{+}$ and $t=\tau,\tau+1,\ldots,T$ is given by
$V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta)=\min_{a\in\{0,1\}}Q^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta,a),$ (15)
@@
$V^{\boldsymbol{\lambda}_{m,T+1:T},\boldsymbol{\mu}_{T+1:T}}_{m,j,T+1}(\delta)=0.$ (16)
@@
$V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta)$
@@
$+\gamma\mathbb{E}\!\left[V^{\boldsymbol{\lambda}_{m,t+1,T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t+1}(\Delta_{m,j}(t+1))\bigg{|}\Delta_{m,j}(t)\!=\!\delta,\!\pi_{m,j}(t)\!=a\!\right].$ (17)
Lemma 1 establishes an optimal decision $\pi^{*}_{m,j}(t)$ for problem (IV-B) by using dynamic programming method. The backward induction method to compute the value function $V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta)$ for $t=\tau,\tau+1,\ldots,T$ is given by
$V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta)$
$=\!w_{m,j}p_{m,j}(\delta)+\!\min_{a\in\{0,1\}}\bigg{\{}\!(1-a)\gamma V^{\boldsymbol{\lambda}_{m,t+1:T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t+1}(\delta+1)$
$+a\left(\lambda_{m,t}+\mu_{t}n_{m,j}+\gamma V^{\boldsymbol{\lambda}_{m,t+1:T},\boldsymbol{\mu}_{t+1:T}}_{m,j,t+1}(1)\right)\bigg{\}}.$ (18)
However, if the AoI $\delta$ can take infinite values, this is computationally intractable. Thus, we restrict the computation of value function to a finite range $\delta=1,2,\ldots,\bar{\delta}$, and approximate $V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\delta)\approx V^{\boldsymbol{\lambda}_{m,t:T},\boldsymbol{\mu}_{t:T}}_{m,j,t}(\bar{\delta})$ for values exceeding this range. In reality, this truncation will have a negligible effect since (i) higher AoI values are rarely visited in practice [51], and (ii) the inference error $p_{m,j}(\delta)$ tends to converge to an upper bound as AoI becomes large, as seen in some recent works [4, 10, 11] and our machine learning experiments in Figs. 7 & 12. The backward induction algorithm has a time complexity of $O(\bar{\delta}T)$.
@@ IV-C Solution to (11)
Next, we solve the dual problem (11). Similar to [13, Proposition 3.1(c)], we can show that the optimal objective value $\bar{p}_{\tau}(\boldsymbol{\lambda}_{\tau:T},\boldsymbol{\mu}_{\tau:T})$ of the problem (IV-A) is concave in $\boldsymbol{\lambda}_{\tau:T}$ and $\boldsymbol{\mu}_{\tau:T}$. Because of the concavity, we can solve the dual problem (11) by the stochastic sub-gradient ascent method. The $(i+1)$-th iteration for the sub-gradient ascent method for all $m=1,2,\ldots$ and $t=\tau,\tau+1,\ldots,T$ is given by
$\lambda_{m,t}(i+1)=$
$\max\bigg{\{}\lambda_{m,t}(i)+\frac{\gamma^{t-\tau}\beta_{m}}{Ki}\left(\sum_{j=1}^{k_{m}}\pi_{m,j}^{*}(t)-C_{m}\right),0\bigg{\}},$ (19)
$\mu_{t}(i+1)=$
$\max\bigg{\{}\mu_{t}(i)+\frac{\gamma^{t-\tau}\beta}{Ki}\left(\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}^{*}(t)n_{m,j}-N\right),0\bigg{\}},$ (20)
where $\beta,\beta_{m}>0$ are the step sizes and $\pi_{m,j}^{*}(t)$ is the optimal action to the $(m,j)$-th sub-problem determined with $\lambda_{m,t}(i),\lambda_{m,t+1}(i),\ldots,\lambda_{m,T}(i)$, and $\mu_{t}(i),\mu_{t+1}(i),\ldots,\mu_{T}(i)$ obtained in the $i$-th iteration.
@@ V Scheduling Policy
V-A Reoptimized Maximum Gain First (MGF) Policy
While the decision $\pi^{*}_{m,j}(t)$ provided in Lemma 1 may violate constraints (5)-(6), we exploit the structure of the decision $\pi^{*}_{m,j}(t)$ to develop a scheduling policy for the original problem (4)-(6). The proposed policy utilizes the notion of gain indices discussed in some recent papers [4, 36, 42]. To determine gain indices for our MTRI problem at time $t$, we use the action value function $Q^{\boldsymbol{\lambda}_{m,t:T}^{*},\boldsymbol{\mu}_{t:T}^{*}}_{m,j,t}(\Delta_{m,j}(t),\pi_{m,j}(t))(\cdot)$ associated with Lagrange multipliers $\boldsymbol{\lambda}_{m,t:T}^{*}$ and $\boldsymbol{\mu}_{t:T}^{*}$.
@@ V-A Reoptimized Maximum Gain First (MGF) Policy
$\alpha_{m,j,t}(\delta)=$ $Q^{\boldsymbol{\lambda}_{m,t:T}^{*},\boldsymbol{\mu}_{t:T}^{*}}_{m,j,t}(\Delta_{m,j}(t),\pi_{m,j}(t))(\delta,0)$
$-Q^{\boldsymbol{\lambda}_{m,t:T}^{*},\boldsymbol{\mu}_{t:T}^{*}}_{m,j,t}(\Delta_{m,j}(t),\pi_{m,j}(t))(\delta,1),$ (21)
where the Lagrange multipliers are obtained after solving (11) with $\tau=t$.